# Benchmarking quantum simulation with neutron-scattering experiments

### 16 Mar 2026 v1

@@ Abstract
A central goal of quantum computation is the realistic simulation of quantum materials. Although quantum processors have advanced rapidly in scale and fidelity, it has remained unclear whether pre–fault-tolerant devices can perform quantitatively reliable material simulations within their limited gate budgets. Here, we demonstrate that a superconducting quantum processor operating on up to 50 qubits can already produce meaningful, quantitative comparisons with inelastic neutron-scattering measurements of KCuF3, a canonical realization of a gapless Luttinger liquid system with a strongly correlated ground state and a spectrum of emergent spinons. The quantum simulation is enabled by a quantum–classical workflow for computing dynamical structure factors (DSFs). The resulting spectra are benchmarked against experimental measurements using multiple metrics, highlighting the impact of circuit depth and circuit fidelity on simulation accuracy. Finally, we extend our simulations to 1D XXZ Heisenberg model with next-nearest neighbor interactions and a strong anisotropy, producing a gapped excitation spectrum, which could be used to describe the CsCoX3 compounds above the Néel temperature. Our results establish a framework for computing DSFs for quantum materials in classically challenging regimes of strong entanglement and long-range interactions, enabling quantum simulations that are directly testable against laboratory measurements.
Recent advances in quantum processors have enabled the investigation of both static  (?, ?) and dynamical  (?, ?, ?) properties through quantum simulations of increasingly large many-body systems. However, even with rapid improvements in scale and fidelity, establishing a clear connection between quantum simulations of materials and corresponding experimental outcomes remains a central challenge. Inelastic neutron scattering (INS) is particularly well suited to bridge the gap between quantum simulation and experiment, as it directly probes the dynamical structure factor (DSF), which can be computed on digital quantum computers from the corresponding two-point correlation functions  (?, ?). Neutrons are a weakly out-of-equilibrium probe and provide a faithful representation of a system’s equilibrium dynamics. The scattered neutrons encode both the energy and momentum of the spin excitations, which can be described by many-body spin Hamiltonians that, in turn, can be directly mapped onto digital quantum computers  (?, ?). In particular, by computing spatiotemporal series of retarded Green’s functions and classically transforming them into momentum and frequency space, quantum processors provide direct access to the correlation functions that underlie renormalization-group descriptions of quantum matter  (?), thereby establishing a connection between microscopic quantum simulation and emergent field theory.
While classical and semiclassical methods exist for simulating INS spectra  (?, ?), their applicability is restricted to regimes where quantum effects are negligible or systems can be approximated by tractable bosonic field theories  (?) (wherein interactions and dynamics can be computed perturbatively about a classical ground state (?)). These approaches become inadequete in strongly interacting systems that exhibit nontrivial collective behavior, or where the quantum ground state does not resemble the classical ground state—such as in quantum spin liquids or at quantum criticality—precisely where the most interesting collective quantum behaviour occurs. More advanced non-perturbative techniques, including exact diagonalization  (?), Krylov-space methods  (?), and tensor-network approaches such as t-DMRG and Matrix Product State (MPS) simulations  (?), can simulate unequal-time correlators and DSFs. However, their performance is constrained by exponential scaling in system size and rapid entanglement growth in long time evolution. Overcoming these classical limitations demands quantum hardware capable of performing coherent evolution over many time steps.
Recent improvements in qubit coherence, gate fidelity, and device stability now support deeper circuit and longer time evolutions, yet questions remain whether it is yet possible to reliably compute the DSF of complex phases on current quantum hardware. If quantum computers can reliably compute DSFs with controlled approximations, the quantum simulated spectra can serve as faithful proxies for INS spectra enabling quantitative comparisons between theoretical models and experimental data. To benchmark the utility of quantum simulations in a setting where cross-validation is possible, we focus on a one-dimensional model that is numerically tractable and thus allows verification of the quantum results. We reproduce the rich dynamics in quasi-1D quantum magnets, including signatures of a 1-D quantum spin liquid described by Luttinger theory, exhibiting fractionalized spinon excitations and a continuum of magnetic scattering   (?, ?, ?). Establishing reliability in this tractable setting is a necessary step toward more general regimes where classical high-performance computing (HPC) methods break down. In those regimes, evaluating the DSF for general local Hamiltonians is believed to be BQP‑hard  (?), and classical HPC approaches may fail to capture the nontrivial collective quantum signatures that quantum processors can efficiently access  (?).
[…]
@@
The spin moments of neutrons are scattered by the those in a magnet, making neutron scattering a staple technique for energy-momentum resolved spectroscopy of quantum magnets. INS using moderated thermal and cold neutrons, typically from a reactor or spallation source, is a non-destructive experimental technique widely used to probe spin dynamics at the atomic scale. When the neutrons interact with a material, they exchange energy and momentum with its microscopic degrees of freedom; therefore, measuring the scattered neutrons provides direct access to the underlying eigenstates and excitation spectra of the material.
Evaluating the INS spectrum involves computing the magnetic neutron scattering cross-section  (?) ($I(q,\omega)$), which is directly related to the DSF $S(q,\omega)$, as follows.
@@
Within the framework of the fluctuation-dissipation theorem  (?), the DSF can be evaluated at experimentally achievable temperatures, assuming there are no phase transitions due to the change in temperature, directly from the imaginary part of the retarded two-point Green’s function, $G^{R}(q,\omega)$, as follows:
@@
When a neutron scatters off a one-dimensional XXZ spin-1/2 chain, it injects a single-spin flip, $S=1$, excitation that fractionalizes into a pair of spinons, as illustrated in Figure 2A. These chargeless quasiparticles carry spin-1/2 and emerge naturally in strongly correlated one-dimensional XXZ spin-1/2 chains. Once created, the two spinons can propagate independently while conserving energy and momentum, giving rise to a broad continuum in the INS spectrum, a signature of fractionalized spin dynamics, in contrast to the conventional sharp magnon modes characteristic of bosonic spin waves. In the XX model, this behavior is clearly reflected in the retarded Green’s function (RGF) which exhibits a light-cone structure (See Figure 2B), consistent with the observations in Ref.  (?). The corresponding DSF displays a continuum (Figure 2C), with spectral weight concentrated near its upper boundary, consistent with theoretical expectations of ballistic spinon propagation in the XX limit.
@@
To quantitatively assess how close the quantum hardware results are to the neutron spectra, we first compute three baseline metrics: the mean squared error (MSE), the Wasserstein distance and the structural similarity index measure (SSIM)  (?). MSE measures the average local pixel-wise deviation between two spectra, while the Wasserstein distance captures their global similarity by evaluating the cost of transporting one spectra into another. SSIM provides a complementary structure-aware comparison; its values range from 1 (perfect similarity) to 0 (no similarity) to -1 (perfect antisimilarity).
MSE, Wasserstein distance, and SSIM all quantitatively confirm that the quantum simulation and the MPS results reproduce the INS spectra with comparable fidelity. As summarized in Figure 2J, the MPS simulation (computed from the DMRG ground state using the same number of qubits, time discretization and discrete Fourier transform procedure) sets an upper bound on the accuracy achievable in the absence of hardware noise. Accordingly, its slightly better MSE and SSIM is expected. Any other visual similarity observed between the quantum simulation and the INS spectra arises from noise-induced data smearing rather than from the underlying dynamics.
Although these baseline metrics provide useful global similarity measures, they remain relatively generic image-based metrics less relevant to condensed-matter applications, such as peak positions, linewidths, spectral-weight distribution, or correlation and entanglement properties. To address these, we turn to additional physics-informed metrics.
In terms of the spectral distribution, the main dispersion carries approximately 48 % of the total spectral weight in the experimental spectrum. The MPS results slightly overestimate this weight at 58.9 %, whereas the ibm_bostonibm\_boston results lie close to the experimental value, yielding weights near 50 %. This apparent agreement, however, is a consequence of noise rather than increased accuracy. Hardware noise reduces the ability to cleanly resolve the main dispersion, effectively smearing the intensity and thus lowering the observable spectral weight relative to MPS result.
@@
As the error rate is reduced from $7.5\times 10^{-3}$ to $1\times 10^{-3}$, we observe substantial improvements in spectrum quality, as reflected in the SSIM, MSE, and nQFI metrics. To further validate the effect of the error rate on agreement with the reference spectrum, we perform quantum experiments on two different generations of IBM’s Heron quantum processors, ibm_kingston and ibm_boston, whose respective error rates for the 50‑qubit experiment are summarized in Figure 3B. Unsurprisingly, the significantly lower gate errors for ibm_boston relative to ibm_kingston lead to a systematically better agreement with the INS spectra, showing lower MSE, higher SSIM and nQFI accuracy. As the median two-qubit error rate of our current device approaches $1\times 10^{-3}$, we are now entering an era in which quantum computers can be used as a tool to simulate INS spectra. Continued advances in error mitigation and suppression techniques will effectively increase the gate budget, enabling the simulation of increasingly complex INS spectra in the near future.
@@
where $J^{\prime}$ and $\epsilon^{\prime}$ are the exchange interaction strength and anisotropy parameter within the NNN pairs. The easy axis (Ising-like) limit of the XXZ Hamiltonian now restricts the anisotropy term such that $0\leq\epsilon\leq 1$ (note that the anisotropy term acts on the X and Y components unlike Equation 3).
@@
Importantly, the weak NNN couplings introduce controlled departures from integrability. This breaks analytical tractability and places the Hamiltonian far from any analytically solvable limit, unlike KCuF3. As a result, the quasi 1D XXZ model provides a regime where direct comparisons between the MPS and quantum simulations of the DSF are particularly meaningful, since an exact analytical benchmark is no longer available.
The quantum simulated DSF is generally consistent with the MPS results. The peak positions of the quantum and MPS results at the $\pi/2$ and $\pi$ points are both within a 10% error relative to the FWHM of the MPS spectrum, as shown by the line scan in Figure 4E. In terms of the entanglement witness, the nQFI of the XX channel at the $\pi$ point is below 1.0 in both the quantum and MPS calculations, indicating a weakly entangled regime with low multipartite entanglement compared to KCuF3, as expected. The corresponding quantum and MPS values are summarized in Figure 4F. Given that the dispersion shapes and excitation energies can be reliably predicted, we are able to qualitatively characterize materials based on the DSF for this class of models; however, experimental verification of this Hamiltonian through INS experiments is still needed.
@@
In this work, we utilize Inelastic Neutron Scattering (INS) spectra as a strategic benchmark for tracking the accelerating progress of quantum simulations. Using KCuF3 and a strongly anisotropic 1D XXZ model with NNN interactions, we show that quantum simulations utilizing up to 50 qubits, serving as proxies for the thermodynamic limit, can faithfully reconstruct the dynamical structure factor of real materials.
Our results reveal that current quantum hardware already captures key emergent quantum phenomena, including the two-spinon continuum present in real materials, as well as how anisotropy and realistic NNN couplings reshape this continuum. This marks a decisive moment: quantum computers are no longer just testbeds; they are emerging as practical tools alongside classical tensor-network methods for studying quantum materials.
Beyond remarkable qualitative agreements, we introduce a set of rigorous quantitative metrics, ranging from standard image-based measures to entanglement-based physical metrics, that remain computable even when classical exact methods struggle. These metrics form a crucial roadmap for monitoring hardware progress. As these figures of merit improve, we anticipate a crossover point where quantum simulation evolves from a benchmarking target into an indispensable tool for condensed matter physicists, enabling discovery in regimes where classical simulation is no longer viable.
The central achievement of this work is establishing a quantum benchmark and demonstrating the utility of current quantum devices. This represents a significant advancement over earlier calculations of DSFs  (?, ?, ?, ?, ?) and paves the way for modeling and understanding systems that are beyond classical tractability. As quantum devices scale to larger lattices, support longer time evolutions, and ultimately handle general two-dimensional non-integrable models, this framework positions quantum simulation as a renormalization-group-compatible engine for constructing complete equilibrium descriptions of quantum materials. The work presented here is the first step along this path, demonstrating that quantum hardware can faithfully reproduce experimentally measured DSF with high accuracy and signaling the beginning of a new era in computational condensed matter physics.
@@ Acknowledgments
The KCuF3 spectrum was previously measured and reported in Refs. (?, ?) using a combination of MAPS (ISIS) and SEQUOIA (ORNL) spectrometers: MAPS for high-energy ($\hbar\omega>8$ meV), SEQUOIA for low-energy ($\hbar\omega<8$ meV).
@@
To prepare the ground-state circuits for the spin models studied in Figure 2 and  4, we perform optimization of Hamiltonian variational ansatz acting on appropriate initial states. The optimization is performed classically, with the loss function chosen to be the fidelity overlap with respect to the MPS ground state computed using DMRG, i.e., $|\langle\psi_{o}|\psi_{\mathrm{GS}}\rangle|^{2}$. For the 1D XX model and the 1D XXZ model at the isotropic point, we adopt the procedure outlined in the work (?): the process begins with pairs of singlets, which form the ground state of a Hamiltonian corresponding to the XXZ model acting only on the odd pairs of the system. We then apply the variational Hamiltonian ansatz on the entire system. This approach enables convergence to the ground state for both cases. For the ground states corresponding to $\epsilon=0.145$, the Néel state is closer to them than the singlet configuration. Therefore, it is more appropriate to start from the Néel state and use a variational Hamiltonian ansatz. Using this approach, we achieve significantly higher fidelities compared to starting from singlet states, reaching approximately $95.9\%$.
@@
For the XX model, we observe that six layers are required to reach a fidelity of approximately $70\%-75\%$. In contrast, for isotropic point, the variational optimization achieves higher fidelities slightly faster, achieving a fidelity close to $85\%$ for the same number of layers. For the simulation results presented in Figure 2, we used five layers. We note that for the simulations shown in Figure 4B, we used the same ground state as in the case without NNN interactions. This choice is justified because these NNN interactions are very small in magnitude, and the fidelity overlap with respect to the true ground state (including NNN interactions) is nearly identical to that obtained without them. The small NNN interaction does not significantly alter the ground-state wavefunction, as the fidelity between the states with and without the NNN interaction is $99.4\%$.

### 25 Mar 2026 v2

@@ Abstract
A central goal of quantum computation is the realistic simulation of quantum materials. Although quantum processors have advanced rapidly in scale and fidelity, it has remained unclear whether pre–fault-tolerant devices can perform quantitatively reliable material simulations within their limited gate budgets. Here, we demonstrate that a superconducting quantum processor operating on up to 50 qubits can already produce meaningful, quantitative comparisons with inelastic neutron-scattering measurements of KCuF3, a canonical realization of a gapless Luttinger liquid system with a strongly correlated ground state and a spectrum of emergent spinons. The quantum simulation is enabled by a quantum–classical workflow for computing dynamical structure factors (DSFs). The resulting spectra are benchmarked against experimental measurements using multiple metrics, highlighting the impact of circuit depth and circuit fidelity on simulation accuracy. Finally, we extend our simulations to a 1D XXZ Heisenberg model with next-nearest-neighbor interactions and a strong anisotropy, producing a gapped excitation spectrum, which could be used to describe the CsCoX3 compounds above the Néel temperature. Our results establish a framework for computing DSFs for quantum materials in classically challenging regimes of strong entanglement and long-range interactions, enabling quantum simulations that are directly testable against laboratory measurements.
Accurately predicting the properties of quantum materials has been enabled by decades of advances in classical computational methods  (?, ?, ?, ?). However, strongly correlated systems with long-range entanglement and complex dynamics remain beyond the reach of these approaches. Quantum computers offer a potential alternative to addressing this challenge  (?), as envisioned by Feynman  (?). Rapid advances in qubit number and gate fidelity have enabled studies of both static  (?, ?) and dynamical  (?, ?, ?) properties of many-body systems at scales beyond exact diagonalization. Yet, despite recent advances in algorithms and hardware, resource estimates suggest that extending these simulations to realistic materials is expected to require circuit depths and error rates that are beyond near-term capabilities  (?). It has therefore been unclear whether current, pre-fault tolerant quantum computers can ever perform quantitatively reliable many-body simulations of quantum materials that can be closely compared with laboratory measurements.
Inelastic neutron scattering (INS) provides an experimentally accessible pathway to bridge quantum simulation with laboratory measurements. In an INS experiment, incident neutrons scatter off magnetic ions in a material, exchanging both energy and momentum with its spins. The measured signal can be used to compute two-point dynamical correlation functions that capture how local excitations propagate through the system in space and time. Expressed in momentum and frequency space, these correlations form the dynamical structure factor (DSF), which directly reveals the spectrum of collective excitations and dispersion relations in a material. As many quantum materials can be modeled by spin-1/2 Hamiltonians that are naturally mapped onto qubit systems, digital quantum processors can compute the DSF  (?, ?, ?, ?, ?) by preparing the ground state, perturbing it locally, evolving under the system Hamiltonian, and measuring two-point correlation functions. This correspondence establishes a direct pathway from quantum circuits to experimentally observable quantities.
Classical approaches to computing DSFs—including semiclassical spin-wave methods  (?), exact diagonalization  (?), Krylov-space techniques  (?), and tensor-network methods such as time-dependent density-matrix renormalization group (t-DMRG)  (?)—have achieved significant success. However, they are fundamentally limited by exponential growth in Hilbert-space dimension or by entanglement growth during real-time evolution, restricting their applicability in strongly interacting regimes such as quantum spin liquids or near quantum criticality. For general local Hamiltonians, evaluating the dynamical structure factor is believed to be BQP-hard, inheriting the computational complexity of simulating real-time quantum dynamics  (?), and highlighting the potential advantage of quantum processors for accessing these observables  (?).
Here we demonstrate that quantum computers can serve as reliable tools for simulating experimentally measurable spectra of quantum materials. We perform a direct experimental validation by benchmarking quantum simulations against INS measurements of the quasi-one-dimensional antiferromagnet KCuF3 , whose spectrum is a canonical realization of a Tomonaga–Luttinger liquid exhibiting fractionalized spinon excitations  (?, ?, ?). Using 50-qubit circuits executed on superconducting quantum processors, we compute spatiotemporal correlation functions and reconstruct the DSF for quantitative comparison with experiment. We show that the agreement between simulation and experiment is enabled by reductions in two-qubit error rates towards 0.1%, and the intrinsic robustness of the DSF estimation to noise-induced damping of the correlation functions. We further extend this framework to a non-integrable XXZ model with next-nearest-neighbor interactions, relevant to CsCoX3 (X = Halides) above their Néel temperature  (?, ?, ?), demonstrating applicability beyond analytically solvable regimes. These results establish a pathway for quantum simulations that are directly testable against laboratory measurements, providing a benchmark for the scientific capabilities of near-term quantum processors.
@@
When the neutrons interact with a material, they exchange energy and momentum with its microscopic degrees of freedom; therefore, measuring the scattered neutrons provides direct access to the underlying eigenstates and excitation spectra of the material. INS using moderated thermal and cold neutrons, typically from a reactor or spallation source, is a non-destructive experimental technique widely used to probe spin dynamics at the atomic scale. Evaluating the INS spectrum involves computing the magnetic neutron scattering cross-section  (?) ($I(q,\omega)$), which is directly related to the DSF $S(q,\omega)$, as follows.
@@
Within the framework of the fluctuation-dissipation theorem  (?), , assuming there are no phase transitions due to the change in temperature, the DSF can be evaluated at experimentally achievable temperatures directly from the imaginary part of the retarded two-point Green’s function, $G^{R}(q,\omega)$, as follows:
@@
When a neutron scatters off a one-dimensional XXZ spin-1/2 chain, it injects a single-spin flip, $S=1$, excitation that fractionalizes into a pair of spinons, as illustrated in Figure 2A. These chargeless quasiparticles carry spin-1/2 and emerge naturally in strongly correlated one-dimensional XXZ spin-1/2 chains. Once created, the two spinons can propagate independently while conserving energy and momentum, giving rise to a broad continuum in the INS spectrum, a signature of fractionalized spin dynamics, in contrast to the conventional sharp magnon modes characteristic of bosonic spin waves. In the XX model, this behavior is clearly reflected in the retarded Green’s function (RGF), which exhibits a light-cone structure (See Figure 2B), consistent with the observations in Ref.  (?). The corresponding DSF displays a continuum (Figure 2C), with spectral weight concentrated near its upper boundary, consistent with theoretical expectations of ballistic spinon propagation in the XX limit.
@@
To quantitatively assess how close the quantum hardware results are to the neutron spectra, we first compute three baseline metrics: the mean squared error (MSE), the Wasserstein distance, and the structural similarity index measure (SSIM)  (?). MSE measures the average local pixel-wise deviation between two spectra, while the Wasserstein distance captures their global similarity by evaluating the cost of transporting one spectrum into another. SSIM provides a complementary structure-aware comparison; its values range from 1 (perfect similarity) to 0 (no similarity) to -1 (perfect antisimilarity).
MSE, Wasserstein distance, and SSIM all quantitatively confirm that the quantum simulation and the MPS results reproduce the INS spectra with comparable fidelity. As summarized in Figure 2J, the MPS simulation (computed from the DMRG ground state using the same number of qubits, time discretization, and discrete Fourier transform procedure) sets an upper bound on the accuracy achievable in the absence of hardware noise. Accordingly, its slightly better MSE and SSIM are expected. Any other visual similarity observed between the quantum simulation and the INS spectra arises from noise-induced data smearing rather than from the underlying dynamics.
Although these baseline metrics provide useful global similarity measures, they remain relatively generic image-based metrics, less relevant to condensed-matter applications, such as peak positions, linewidths, spectral-weight distribution, or correlation and entanglement properties. To address these, we turn to additional physics-informed metrics.
In terms of the spectral distribution, the main dispersion carries approximately 48 % of the total spectral weight in the experimental spectrum. The MPS results slightly overestimate this weight at 58.9 %, whereas the ibm_bostonibm\_boston results lie close to the experimental value, yielding weights near 50 %. This apparent agreement, however, is a consequence of noise rather than an increase in accuracy. Hardware noise reduces the ability to cleanly resolve the main dispersion, effectively smearing the intensity and thus lowering the observable spectral weight relative to the MPS result.
@@
As the error rate is reduced from $7.5\times 10^{-3}$ to $1\times 10^{-3}$, we observe substantial improvements in spectrum quality, as reflected in the SSIM, MSE, and nQFI metrics. To further validate the effect of the error rate on agreement with the reference spectrum, we perform quantum experiments on two different generations of IBM’s Heron quantum processors, ibm_kingston and ibm_boston, whose respective error rates for the 50‑qubit experiment are summarized in Figure 3B. Unsurprisingly, the significantly lower gate errors for ibm_boston relative to ibm_kingston lead to a systematically better agreement with the INS spectra, showing lower MSE, higher SSIM, and nQFI accuracy. As the median two-qubit error rate of our current device approaches $1\times 10^{-3}$, we are now entering an era in which quantum computers can be used as a tool to simulate INS spectra. Continued advances in error mitigation and suppression techniques will effectively increase the gate budget, enabling the simulation of increasingly complex INS spectra in the near future.
@@
where $J^{\prime}$ and $\epsilon^{\prime}$ are the exchange interaction strength and anisotropy parameter within the NNN pairs. The easy axis (Ising-like) limit of the XXZ Hamiltonian now restricts the anisotropy term such that $0\leq\epsilon\leq 1$ (note that the anisotropy term acts on the X and Y components, unlike Equation 3).
@@
Importantly, the weak NNN couplings introduce controlled departures from integrability. This breaks analytical tractability and places the Hamiltonian far from any analytically solvable limit, unlike KCuF3. As a result, the quasi-1D XXZ model provides a regime where direct comparisons between the MPS and quantum simulations of the DSF are particularly meaningful, since an exact analytical benchmark is no longer available.
The quantum-simulated DSF is generally consistent with the MPS results. The peak positions of the quantum and MPS results at the $\pi/2$ and $\pi$ points are both within a 10% error relative to the FWHM of the MPS spectrum, as shown by the line scan in Figure 4E. In terms of the entanglement witness, the nQFI of the XX channel at the $\pi$ point is below 1.0 in both the quantum and MPS calculations, indicating a weakly entangled regime with low multipartite entanglement compared to KCuF3, as expected. The corresponding quantum and MPS values are summarized in Figure 4F. Given that the dispersion shapes and excitation energies can be reliably predicted, we are able to qualitatively characterize materials based on the DSF for this class of models; however, experimental verification of this Hamiltonian through INS experiments is still needed.
@@
In this work, we demonstrate that quantum simulation of real materials on pre‑fault‑tolerant, programmable quantum hardware is no longer an elusive goal. Using a superconducting‑qubit quantum processor, we reconstruct the dynamical structure factor of KCuF3, a quasi‑one‑dimensional Luttinger liquid, and a strongly anisotropic one‑dimensional XXZ model with next‑nearest‑neighbor interactions, which is used to describe CsCoX3 above Néel temperature. For KCuF3, the simulated spectra are directly compared with inelastic neutron scattering measurements performed at 6 K. These quantum simulations faithfully capture key emergent quantum phenomena observed in real materials, including the two‑spinon continuum and the manner in which anisotropy and realistic next‑nearest‑neighbor couplings reshape this continuum. Together, these results establish that quantum computers are moving beyond proof‑of‑principle testbeds and are beginning to function as practical scientific tools for the study of quantum materials.
Beyond qualitative agreement, we introduce a suite of quantitative performance metrics—ranging from standard image‑based measures to physically motivated entanglement‑based diagnostics—that remain computable even in regimes where exact classical methods become intractable. This represents a substantial advance over prior quantum calculations of dynamical structure factors  (?, ?, ?, ?, ?) and enables principled assessment of quantum simulations in increasingly complex settings. These metrics provide a roadmap for tracking progress in quantum hardware capabilities. As these figures of merit improve, we anticipate a crossover in which quantum simulation transitions from a benchmarking exercise to an indispensable tool for condensed‑matter physics, enabling exploration of regimes that lie beyond the reach of classical simulation. Looking ahead, as quantum devices scale to larger lattices, support longer time evolutions, and address generic two‑dimensional non‑integrable models, this framework positions quantum simulation as an engine for constructing comprehensive equilibrium descriptions of quantum materials.
@@ Acknowledgments
The KCuF3 spectrum was previously measured and reported in Ref. (?) using the MAPS spectrometer (ISIS) spectrometer. For this experiment the $c$-axis was oriented in the scattering plane perpendicular to the incident beam. Data were corrected for the Cu2+ form factor and integrated over all directions perpendicular to the chains. Because KCuF3 has weak magnetic order below its ordering temperature $T_{N}=39$ K, there are subtle differences between the longitudinal and transverse responses at low energy $\tilde{\omega}<0.5$ (?), but these are minor effects and the spectrum overall matches very closely to the perfect 1D isotropic model (?).
@@
To prepare the ground-state circuits for the spin models studied in Figures 2 and  4, we perform optimization of Hamiltonian variational ansatz acting on appropriate initial states. The optimization is performed classically, with the loss function chosen to be the fidelity overlap with respect to the MPS ground state computed using DMRG, i.e., $|\langle\psi_{o}|\psi_{\mathrm{GS}}\rangle|^{2}$. For the 1D XX model and the 1D XXZ model at the isotropic point, we adopt the procedure outlined in the work (?): the process begins with pairs of singlets, which form the ground state of a Hamiltonian corresponding to the XXZ model acting only on the odd pairs of the system. We then apply the variational Hamiltonian ansatz on the entire system. This approach enables convergence to the ground state for both cases. For the ground states corresponding to $\epsilon=0.145$, the Néel state is closer to them than the singlet configuration. Therefore, it is more appropriate to start from the Néel state and use a variational Hamiltonian ansatz. Using this approach, we achieve significantly higher fidelities compared to starting from singlet states, reaching approximately $95.9\%$.
@@
Although the fidelity decreases with system size, this behavior is expected because both the XX and XXZ models lie in critical phases in the thermodynamic limit. However, for finite systems up to 50 qubits, as considered in this work, even with just six ansatz layers, we obtain fidelities of approximately $70\%-75\%$ for the XX model and close to $85\%$ at the isotropic point of the XXZ model. These fidelities are sufficient to accurately capture the growth of the retarded Green’s function over the simulated time window and to reproduce the corresponding DSF spectra. This further highlights that reliable DSF simulation does not require systems much larger than $\sim 100$ qubits.
For the simulation results presented in Figure 2, we used five layers. We note that for the simulations shown in Figure 4B, we used the same ground state as in the case without NNN interactions. This choice is justified because these NNN interactions are very small in magnitude, and the fidelity overlap with respect to the true ground state (including NNN interactions) is nearly identical to that obtained without them. The small NNN interaction does not significantly alter the ground-state wavefunction, as the fidelity between the states with and without the NNN interaction is $99.4\%$.
@@
Spectrum Broadening
For experimental measurements, spectral features broaden at finite temperature because thermal fluctuations increase the number of accessible transitions and reduce quasiparticle lifetimes. In the quantum simulation results, noise generally damps the measured observables, causing a broadening of the signal.
Consistent with these effects, the full width at half maximum (FWHM) of the KCuF3 spectrum at the $\pi/2$ point is larger at higher temperatures and higher error rates, as shown in Figure S10. Because quantum noise effectively increases the space of measurement outcomes, it can be interpreted as mimicking certain finite‑temperature effects, suggesting the possibility of defining an effective temperature associated with the noise. However, establishing such a correspondence lies beyond the scope of the present work.
@@
Figure S10: Comparison of the KCuF3 spectrum under finite-temperature conditions and depolarization noise. The spectrum at 75K is replotted based on the data in Ref.  (?). The full width at half maximum increases as the measurement temperature and error rate increase. The line scan is measured at the $\pi/2$ point.